# The adelic closure of triangle groups

### v1

@@ 3. The $2$-adic image
where the first $\mathbf{F}_{2}$ is generated by the image of the matrices ±$\pm I$. (As noted in [McM24, §1.V], the choice of signs in the definition of $\Delta$ ensures that $-I\in\Delta$.) Because $\Delta(2)$ is normal in $\Delta$, and since this subspace is invariant under conjugation by $\Delta/\Delta(2)$, it suffices to show that a set of normal generators for $\Delta(2)$ maps to this space. But we know that, with $d=n/2$ if $n$ is even and $n$ otherwise, that
$\Delta/\Delta(2)\simeq\langle T,U|T^{2},U^{2},R^{d},R=(UT)\rangle.$
So it suffices to check the claim for $T^{2}$, $U^{2}$, and $R^{d}$. The image of the latter is given by $-I$. On the other hand, we find that (modulo $4$):
@@
Since we have already seen that the image of $\Delta(2)/\Delta(4)$ in $\mathcal{O}/2\oplus\widetilde{Q}$ lands in $\mathbf{F}_{2}\oplus\widetilde{Q}$ and contains the copy of $\mathbf{F}_{2}$ generated by $-I$, this certainly shows that the image is the entire space $\mathbf{F}_{2}\oplus\widetilde{Q}$. Thus it suffices to prove that $\widehat{\Delta}_{2}(4)=\widehat{\Gamma}_{2}(4)$, or that the image of $\Delta(2^{r})$ surjects onto $\Gamma(2^{r})/\Gamma(2^{r+1})$ for every $r\geq 2$. We prove this by induction. Note that
@@
Suppose either that $r\geq 2$ and that the image is surjective, or that $r=1$ and (by assumption) the image contains $\widetilde{Q}$. In both cases, the image certainly contains $\widetilde{Q}$. Hence, any pair of elements in $M_{0}(\mathcal{O}/2)$, there exist elements
@@
$[I+2^{r}A\bmod 2^{r+1},I+2B\bmod 4]=I+2^{r+1}(AB-BA)\bmod 2^{r+2}$
Now replacing $A$ and $B$ by $A+\lambda I$ and $B+\mu I$ does’t change $AB-BA$. On the other hand, the image of the map
@@
$X_{i,j}:=(R^{i}T)^{2}(R^{j}T)^{-2}=I+\frac{2}{\zeta-1}E\left(\begin{matrix}0&\zeta(\zeta^{-2i-1}-\zeta^{-2j-1})\\ \zeta^{2i+1}-\zeta^{2j+1}&0\end{matrix}\right)E^{-1}\mod 4.$
@@
with $i$ and $j$ both odd. Now let $\mathfrak{p}|2$ in $\mathcal{O}_{K}$, and let $\mathfrak{p}$ be a prime in $\mathcal{O}_{L}$ above $\mathfrak{p}$. Note that there are at most two primes above $\mathfrak{p}$ in $\mathcal{O}_{L}$, and the other has the form $c\mathfrak{p}$ where $c$ is given by complex conjugation. By Lemma 2.6, we may, depending on the parity of $r$, find an $B$ such that the lower left corner has any given valuation at $\mathfrak{p}$ and higher valuation at all other primes. Moreover, since the upper right corner is obtained by acting by complex conjugation, it has the same valuation at $c\mathfrak{p}$ and higher valuation at all other primes. Reverting back to the basis over $\mathcal{O}_{K}$ and to the matrix $A$, the fact that $E$ is $2$-adically invertible means that since all the entries in the original matrix were trivial modulo $P^{r}\mathfrak{p}^{\prime}$ for every prime $\mathfrak{p}^{\prime}$ except $\mathfrak{p}$ means that $A\in\Gamma(P^{r}\cdot P/\mathfrak{p})$ and moreover the image of $A$ in $\Gamma(P^{r})/\Gamma(P^{r+1})$ is non-zero. Moreover, it is non-scalar as well, since otherwise (conjugating again by $E$) it would still be scalar, and by construction it is not of this form. It follows that its image in this quotient lies in precisely one copy of $Q$ associated to the factor $\Gamma(P^{r})/\Gamma(P^{r}\mathfrak{p})$. Hence, because $Q$ is irreducible over $\mathbf{F}_{2}$, we deduce that the image of $\Gamma(P^{r})$ contains all $d$ copies of $Q$, as desired. ∎
@@ 4. The adelic Image
Since $\zeta^{p^{n}}\equiv\eta\bmod\mathfrak{p}$, the images of $\Delta_{n}$ and $\Delta_{m}$ coincide (up to conjugation in $\mathrm{Gal}(K/\mathbf{Q})$ by the inverse of $[p^{k}]\in(\mathbf{Z}/m\mathbf{Z})^{\times}/\pm 1$) after the targets are identified by equation (10). This reduces the claim for $\Delta_{n}$ to the for $\Delta_{m}$, and thus we may assume that $(n,p)=1$.
By [Rib75, Lemma 3.3], it suffices to prove the image surjects onto any pair of factors (note that we are in the situation where we may assume that $|\mathbf{F}|\geq 5$ so $\mathrm{SL}_{2}(\mathbf{F})$ is perfect.) By Goursat’s lemma, the result follows unless the image is a graph $(x,\phi(x))$ for some isomorphism $\phi:\mathrm{SL}_{2}(\mathbf{F})\rightarrow\mathrm{SL}_{2}(\mathbf{F})$. It follows by considering the image of $U$ and $T$ that $\phi$ fixes
@@
Let us examine the image of $T$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$ and $\mathrm{SL}_{2}(\mathcal{O}/\sigma\mathfrak{p})$ respectively for a pair of primes $\sigma\mathfrak{p}\neq\mathfrak{p}$. We fix an isomorphism $\mathbf{F}=\mathcal{O}/\mathfrak{p}$, so that the image of $\zeta+\zeta^{-1}$ in $\mathcal{O}/\mathfrak{p}$ is $\zeta+\zeta^{-1}$. The image of $\zeta+\zeta^{-1}$ in $\mathbf{F}=\mathcal{O}/\sigma\mathfrak{p}$ has the form $\sigma^{-1}(\zeta+\zeta^{-1})$ in $\mathcal{O}/\mathfrak{p}$. The assumption that $\mathfrak{p}\neq\sigma\mathfrak{p}$ is the assumption that $\sigma$ is not in the subgroup of $\mathrm{Gal}(K/\mathbf{Q})$ generated by the Frobenius element $\mathrm{Frob}_{p}\in\mathrm{Gal}(K/\mathbf{Q})$ (which would mean that $\mathfrak{p}=\sigma\mathfrak{p}$). All automorphisms of $\mathrm{SL}_{2}(\mathbf{F})$ come from diagonal automorphisms (conjugation by $\mathrm{GL}_{2}(\mathbf{F})$) and the action of $\mathrm{Gal}(\mathbf{F}/\mathbf{F}_{p})$, which is generated by Frobenius. The latter fixes $U$ and preserves the subspace of upper triangular unipotent matrices. The centralizer of $U$ in $\mathrm{GL}_{2}(\mathbf{F})$ meanwhile consists of the group generated by scalar matrices and lower triangular unipotent matrices. The only such automorphisms which send an upper triangular unipotent matrix to another upper triangular unipotent matrix are trivial. Thus $\phi$ has to be a Galois automorphism, and thus it must be the case that:
@@
for some $i$ where $\sigma\in\mathrm{Gal}(K/\mathbf{Q})$ is not in the group generated by $\mathrm{Frob}_{p}$. This is impossible if the minimal polynomial of $\zeta+\zeta^{-1}$ is separated modulo $p$. But the ring of integers in $\mathbf{Q}(\zeta+\zeta^{-1})$ is $\mathbf{Z}[\zeta+\zeta^{-1}]$, and since we are assuming that $\zeta$ has order $(m,p)=1$, this field is unramified at $p$, and thus we are done. ∎
@@ 4.4. Ramified Primes
We now deal with (odd) ramified primes. As we have seen, if $n=2p^{k}$ and $k\geq 1$, then the mod-$\mathfrak{p}$ representations themselves only have cyclic image, so we defer this case until Section 4.7. We begin with some preliminaries. or any integer $d\geq 2$, we certainly have an exact sequence of groups:
(11) $0\rightarrow M\rightarrow\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})\rightarrow\mathrm{SL}_{2}(\mathbf{F})\rightarrow 0,$
@@
$[I+\pi^{i}A,1+\pi^{j}B]=1+\pi^{i+j}(AB-BA)\bmod\pi^{i+j+1}.$
But since $AB-BA$ generates $M^{0}(\mathbf{F})$ (assuming as we are that $p>2$), we see that $[M,M]$ inductively contains $M(\mathfrak{p}^{i})$ starting with $i=d$ and descending down to $i=2$. Hence $\Phi(M)\supseteq[M,M]$ contains this subgroup as well. But $M/M(\mathfrak{p}^{2})$ is elementary $p$-abelian, so $\Phi(M)\subseteq M(\mathfrak{p}^{2})$ and thus $\Phi(M)=M(\mathfrak{p}^{2})$. ∎
@@
Let us consider the image $G$ of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})$ for some integer $d\geq 2$, and let $G(\mathfrak{p})$ be the kernel of the map $G\rightarrow\mathrm{SL}_{2}(\mathbf{F})$. We know, by assumption, that the image of $G$ in $\mathrm{SL}_{2}(\mathbf{F})$ is either everything or has index $6$ in the one exceptional case. But in either case, the adjoint representation is irreducible as an $\mathbf{F}_{p}[\mathrm{SL}_{2}(\mathbf{F})]$-module. From Lemma 4.5, there is a map:
@@
If $G(\mathfrak{p})$ surjects onto $M/\Phi(M)$ then $G\simeq M$ by Burnside’s basis theorem [Hal59, Thm 12.2.1]. Since $d\geq 2$ was arbitrary, the result follows. Hence it suffices to show that $G(\mathfrak{p})/G(\mathfrak{p}^{2})\subseteq\Gamma(\mathfrak{p})/\Gamma(\mathfrak{p}^{2})$ is non-zero, since the adjoint representation is irreducible. In particular, it suffices to write down a single element with non-zero image. Write $n=mp^{k}$ with $(m,p)=1$ and $m\neq 2$. Let $v$ denote the valuation on $K_{\mathfrak{p}}$ so that $v(\mathfrak{p}\setminus\mathfrak{p}^{2})=1$. Note that $v$ extends to any finite extension of $K_{\mathfrak{p}}$. If $\gamma\in G(\mathfrak{p}^{2})$, then any eigenvalue $\lambda\in\overline{K}$ of $\gamma$ satisfies
@@
Thus it suffices to write down an element $\gamma$ which is trivial modulo $\mathfrak{p}$ (so $\gamma\in G(\mathfrak{p})$) and has eigenvalues which do not satisfy (12) so $\gamma\notin G(\mathfrak{p}^{2})$. The element $R$ has eigenvalues $\zeta$ and $\zeta^{-1}$. If $n=p^{k}m$ with $m>2$, then $\zeta$ and $\zeta^{-1}$ are distinct modulo $\mathfrak{p}$, and thus the reduction of $R$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$ is semisimple. It follows that $\gamma=R^{m}$ is trivial modulo $\mathfrak{p}$ and has eigenvalues $\zeta^{m}=\zeta_{p^{k}}$ and $\zeta^{-1}_{p^{k}}$. If $m>2$, then $\mathbf{Q}(\zeta)/\mathbf{Q}(\zeta+\zeta^{-1})$ is unramified at $p$, and thus $v(\zeta_{p^{k}}-1)=1$. Thus $\gamma$ gives the desired element, and so we are done unless $m\leq 2$. We are assuming that $m\neq 2$. If $m=1$, then the reduction of $R\bmod\mathfrak{p}$ is not semisimple, and hence some alternate argument is required. We now assume that $n=p^{k}$. There is now a unique prime $\pi=\zeta+\zeta^{-1}-2$ above $p$ in $K$ with residue field $\mathbf{F}_{p}$. We use the following ad hoc construction. Let $V=RU=UTU$; we will now show that $\gamma=V^{p^{2}-1}$ lies in $G(\mathfrak{p})$ and is non-trivial in $G(\mathfrak{p})/G(\mathfrak{p}^{2})$. We first note that the mod $\mathfrak{p}$ reduction of $V$ is as follows:
@@
The Newton polygon of this polynomial has slopes 00 and $1$, from which we deduce that $v(\eta/\epsilon-1)=1$, as desired. ∎

### v2

@@ 3. The $2$-adic image
where the first $\mathbf{F}_{2}$ is generated by the image of the matrices $\pm I$. (As noted in [McM24, §1.V], the choice of signs in the definition of $\Delta$ ensures that $I\in\Delta$.) Because $\Delta(2)$ is normal in $\Delta$, and since this subspace is invariant under conjugation by $\Delta/\Delta(2)$, it suffices to show that a set of normal generators for $\Delta(2)$ maps to this space. But we know that, with $d=n/2$ if $n$ is even and $n$ otherwise, that
$\Delta/\Delta(2)\simeq\langle T,U|T^{2},U^{2},R^{d},R=UT\rangle.$
So it suffices to check the claim for $T^{2}$, $U^{2}$, and $R^{d}$. The image of the latter is given by I-I. On the other hand, we find that (modulo $4$):
@@
Since we have already seen that the image of $\Delta(2)/\Delta(4)$ in $\mathcal{O}/2\oplus\widetilde{Q}$ lands in $\mathbf{F}_{2}\oplus\widetilde{Q}$ and contains the copy of $\mathbf{F}_{2}$ generated by I-I, this certainly shows that the image is the entire space $\mathbf{F}_{2}\oplus\widetilde{Q}$. Thus it suffices to prove that $\widehat{\Delta}_{2}(4)=\widehat{\Gamma}_{2}(4)$, or that the image of $\Delta(2^{r})$ surjects onto $\Gamma(2^{r})/\Gamma(2^{r+1})$ for every $r\geq 2$. We prove this by induction. Note that
@@
Suppose either that $r\geq 2$ and that the image is surjective, or that $r=1$ and (by assumption) the image contains $\widetilde{Q}$. In both cases, the image certainly contains $\widetilde{Q}$. Hence for any pair of elements in $M_{0}(\mathcal{O}/2)$, there exist elements
@@
$I+2^{r}A\bmod 2^{r+1},I+2B\bmod 4]=I+2^{r+1}(AB-BA)\bmod 2^{r+2}$
Now replacing $A$ and $B$ by $A+\lambda I$ and $B+\mu I$ does’t change ABBAAB-BA. On the other hand, the image of the map
@@
$X_{i,j}:=(R^{i}T)^{2}(R^{j}T)^{-2}=I+\frac{2}{\zeta-1}E\left(\begin{matrix}0&\zeta(\zeta^{-2i-1}-\zeta^{-2j-1})\\ \zeta^{2i+1}-\zeta^{2j+1}&0\end{matrix}\right)E^{-1}\mod 4,$
@@
with $i$ and $j$ both odd.
Now let $\mathfrak{p}|2$ in $\mathcal{O}_{K}$, and let $\mathfrak{p}$ be a prime in $\mathcal{O}_{L}$ above $\mathfrak{p}$. Note that there are at most two primes above $\mathfrak{p}$ in $\mathcal{O}_{L}$, and the other has the form $c\mathfrak{p}$ where $c$ is given by complex conjugation. By Lemma 2.6, we may, depending on the parity of $r$, find an $B$ such that the lower left corner has any given valuation at $\mathfrak{p}$ and higher valuation at all other primes. Moreover, since the upper right corner is obtained by acting by complex conjugation, it has the same valuation at $c\mathfrak{p}$ and higher valuation at all other primes. Reverting back to the basis over $\mathcal{O}_{K}$ and to the matrix $A$, the fact that $E$ is $2$-adically invertible means that since all the entries in the original matrix were trivial modulo $P^{r}\mathfrak{p}^{\prime}$ for every prime $\mathfrak{p}^{\prime}$ except $\mathfrak{p}$ means that $A\in\Gamma(P^{r}\cdot P/\mathfrak{p})$ and moreover the image of $A$ in $\Gamma(P^{r})/\Gamma(P^{r+1})$ is non-zero. Moreover, it is non-scalar as well, since otherwise (conjugating again by $E$) it would still be scalar, and by construction it is not of this form. It follows that its image in this quotient lies in precisely one copy of $Q$ associated to the factor $\Gamma(P^{r})/\Gamma(P^{r}\mathfrak{p})$. Hence, because $Q$ is irreducible over $\mathbf{F}_{2}$, we deduce that the image of $\Gamma(P^{r})$ contains all $d$ copies of $Q$, as desired. ∎
@@ 4. The adelic image
Since $\zeta^{p^{n}}\equiv\eta\bmod\mathfrak{p}$, the images of $\Delta_{n}$ and $\Delta_{m}$ coincide (up to conjugation in $\mathrm{Gal}(K/\mathbf{Q})$ by the inverse of $p^{k}]\in(\mathbf{Z}/m\mathbf{Z})^{\times}/\pm 1$) after the targets are identified by equation (10). This reduces the claim for $\Delta_{n}$ to the for $\Delta_{m}$, and thus we may assume that $(n,p)=1$.
By [Rib75, Lemma 3.3], it suffices to prove the image surjects onto any pair of factors (note that we are in the situation where we may assume that $|\mathbf{F}|\geq 5$ so $\mathrm{SL}_{2}(\mathbf{F})$ is perfect). By Goursat’s lemma, the result follows unless the image is a graph $(x,\phi(x))$ for some isomorphism $\phi:\mathrm{SL}_{2}(\mathbf{F})\rightarrow\mathrm{SL}_{2}(\mathbf{F})$. It follows by considering the image of $U$ that $\phi$ fixes
@@
Let us examine the image of $T$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$ and $\mathrm{SL}_{2}(\mathcal{O}/\sigma\mathfrak{p})$ respectively for a pair of primes $\sigma\mathfrak{p}\neq\mathfrak{p}$. We fix an isomorphism $\mathbf{F}=\mathcal{O}/\mathfrak{p}$, so that the image of $\zeta+\zeta^{-1}$ in $\mathcal{O}/\mathfrak{p}$ is $\zeta+\zeta^{-1}$. The image of $\zeta+\zeta^{-1}$ in $\mathbf{F}=\mathcal{O}/\sigma\mathfrak{p}$ has the form $\sigma^{-1}(\zeta+\zeta^{-1})$ in $\mathcal{O}/\mathfrak{p}$. Thus
(11) $\phi\left(\left(\begin{matrix}1&0\\ 2&1\end{matrix}\right)\right)$ $=\left(\begin{matrix}1&0\\ 2&1\end{matrix}\right),$
$\phi\left(\left(\begin{matrix}1&2+\zeta+\zeta^{-1}\\ 0&1\end{matrix}\right)\right)$ $=\left(\begin{matrix}1&2+\sigma^{-1}(\zeta+\zeta^{-1})\\ 0&1\end{matrix}\right).$
The assumption that $\mathfrak{p}\neq\sigma\mathfrak{p}$ is the assumption that $\sigma$ is not in the subgroup of $\mathrm{Gal}(K/\mathbf{Q})$ generated by the Frobenius element $\mathrm{Frob}_{p}\in\mathrm{Gal}(K/\mathbf{Q})$ (which would mean that $\mathfrak{p}=\sigma\mathfrak{p}$). All automorphisms of $\mathrm{SL}_{2}(\mathbf{F})$ can be written as a product of a diagonal automorphism (conjugation by $\mathrm{GL}_{2}(\mathbf{F})$) and the action of $\mathrm{Gal}(\mathbf{F}/\mathbf{F}_{p})$, which is generated by Frobenius. (The latter are called field automorphisms or Galois automorphisms.) The centralizer of a unipotent element in $\mathrm{GL}_{2}(\mathbf{F})$ meanwhile consists of the group generated by that element together with scalar matrices (and the latter give trivial automorphisms.) In particular, the only $\phi$ compatible with the first line of (11) is a product of a field automorphism with conjugation by some power of the given element. The only such automorphisms which send a (non-diagonal) upper triangular unipotent matrix to another upper triangular unipotent matrix are the field automorphisms. Thus $\phi$ has to be a Galois automorphism, and thus it must be the case that:
@@
for some $i$ where $\sigma\in\mathrm{Gal}(K/\mathbf{Q})$ is not in the group generated by $\mathrm{Frob}_{p}$. In particular, these are different global conjugates of $\zeta+\zeta^{-1}$ and hence such a congruence is impossible if the minimal polynomial of $\zeta+\zeta^{-1}$ is separated modulo $p$. But the ring of integers in $\mathbf{Q}(\zeta+\zeta^{-1})$ is $\mathbf{Z}[\zeta+\zeta^{-1}]$, and since we are assuming that $\zeta$ has order $(m,p)=1$, this field is unramified at $p$, and thus we are done. ∎
@@ 4.4. Ramified Primes
We now deal with (odd) ramified primes. As we have seen, if $n=2p^{k}$ and $k\geq 1$, then the mod-$\mathfrak{p}$ representations themselves only have cyclic image, so we defer this case until Section 4.7. We begin with some preliminaries. For any integer $d\geq 2$, we certainly have an exact sequence of groups:
(12) $0\rightarrow M\rightarrow\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})\rightarrow\mathrm{SL}_{2}(\mathbf{F})\rightarrow 0.$
@@
$I+\pi^{i}A,1+\pi^{j}B]=1+\pi^{i+j}(AB-BA)\bmod\pi^{i+j+1}.$
But since ABBAAB-BA generates $M^{0}(\mathbf{F})$ (assuming as we are that $p>2$), we see that $[M,M]$ inductively contains $M(\mathfrak{p}^{i})$ starting with $i=d$ and descending down to $i=2$. Hence $\Phi(M)\supseteq[M,M]$ contains this subgroup as well. But $M/M(\mathfrak{p}^{2})$ is elementary $p$-abelian, so $\Phi(M)\subseteq M(\mathfrak{p}^{2})$ and thus $\Phi(M)=M(\mathfrak{p}^{2})$. ∎
@@
Let us consider the image $G$ of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})$ for some integer $d\geq 2$, and let $G(\mathfrak{p})$ be the kernel of the map $G\rightarrow\mathrm{SL}_{2}(\mathbf{F})$. We know, by assumption, that the image of $G$ in $\mathrm{SL}_{2}(\mathbf{F})$ is either everything or has index $6$ in the one exceptional case. But in either case, the adjoint representation is irreducible as an $\mathbf{F}_{p}[\mathrm{SL}_{2}(\mathbf{F})]$-module.
From Lemma 4.5, there is a map:
@@
If $G(\mathfrak{p})$ surjects onto $M/\Phi(M)$ then $G\simeq M$ by Burnside’s basis theorem [Hal59, Thm 12.2.1]. Since $d\geq 2$ was arbitrary, the result follows. Hence it suffices to show that
$G(\mathfrak{p})/G(\mathfrak{p}^{2})\subseteq\Gamma(\mathfrak{p})/\Gamma(\mathfrak{p}^{2})$
is non-zero, since the adjoint representation is irreducible. In particular, it suffices to write down a single element with non-zero image.
Write $n=mp^{k}$ with $(m,p)=1$ and $m\neq 2$. Let $v$ denote the valuation on $K_{\mathfrak{p}}$ so that $v(\mathfrak{p}\setminus\mathfrak{p}^{2})=1$. Note that $v$ extends to any finite extension of $K_{\mathfrak{p}}$. If $\gamma\in G(\mathfrak{p}^{2})$, then any eigenvalue $\lambda\in\overline{K}$ of $\gamma$ satisfies
@@
Thus it suffices to write down an element $\gamma$ which is trivial modulo $\mathfrak{p}$ (so $\gamma\in G(\mathfrak{p})$) and has eigenvalues which do not satisfy (13) so $\gamma\notin G(\mathfrak{p}^{2})$. The element $R$ has eigenvalues $\zeta$ and $\zeta^{-1}$. If $n=p^{k}m$ with $m>2$, then $\zeta$ and $\zeta^{-1}$ are distinct modulo $\mathfrak{p}$, and thus the reduction of $R$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$ is semisimple. It follows that $\gamma=R^{m}$ is trivial modulo $\mathfrak{p}$ and has eigenvalues $\zeta^{m}=\zeta_{p^{k}}$ and $\zeta^{-1}_{p^{k}}$. If $m>2$, then $\mathbf{Q}(\zeta)/\mathbf{Q}(\zeta+\zeta^{-1})$ is unramified at $p$, and thus $v(\zeta_{p^{k}}-1)=1$. Thus $\gamma$ gives the desired element, and so we are done unless $m\leq 2$. We are assuming that $m\neq 2$. If $m=1$, then the reduction of $R\bmod\mathfrak{p}$ is not semisimple, and hence some alternate argument is required, which we turn to now.
We now assume that $m=1$ so $n=p^{k}$. There is now a unique prime $\pi=\zeta+\zeta^{-1}-2$ above $p$ in $K$ with residue field $\mathbf{F}_{p}$. We use the following ad hoc construction. Let $V=RU=UTU$; we will now show that $\gamma=V^{p^{2}-1}$ lies in $G(\mathfrak{p})$ and is non-trivial in $G(\mathfrak{p})/G(\mathfrak{p}^{2})$. We first note that the mod $\mathfrak{p}$ reduction of $V$ is as follows:
@@
The Newton polygon of this polynomial has slopes 0 and $1$, from which we deduce that $v(\eta/\epsilon-1)=1$, as desired. ∎